Startup Night 5 - 20150907 UT

Observers: JHill, (LBTO) BRothberg, OKuhn (Tucson)
Telescope Operator: DHuerta (LBTO)
AO Support:
Telescope Support: MWagner (LBTO)
Software Support: CBiddick
Instrument Support: JMorris (LBTO)
Engineering Support: DAshby

Summary

Finally the rain has stopped. We got to test the right shutter door and adjust the open limit switches, but nothing else.

Preparations

James H, Mike W, Jared and Chris spent all day getting the elevation axis pads adjusted with low friction.

John H, Mike W and Dan C adjusted the open limit flags on the Left Shutter Door. They are working, but both flags need to be rebuilt.

Dan C worked on the Vent Door software, and continues to do that as we observe.

We did NOT get time to revive the AGW2 guide camera today.

John H added JKerns alignment of M3 for RFBG at -25.727 deg into positions.conf for all the right focal stations. Compare the previous setting of 25.830 deg.

Details

01:00 Opened before sunset for shutter door tests. Sky is spectroscopically clear with wisps of low clouds/fog blowing around.

LBC Wake-up

Close

02:52 Close for heavy clouds. Just as the swingarm reconfiguration to LBC was finishing. Adjusted Left Lower Close limit switch on the shutter door.

-- %USERSIG{JohnHill - 2015-09-07}%
Topic revision: r1 - 07 Sep 2015, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback